A sphere of radius R contains space charge density ρ(r)=A(R-r) for 0

Express A in terms of Q and R.


Find the electric field inside and outside the sphere

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
volume element, dv = 4*pi*r2*dr
integration between 0 to R , ρ(r)dv =Q
Integration [4pi*r2*dr A(R-r)]= Q
Integration [4pi*r2*dr A(R-r)]= Q/4pi
AIntegration [r2R-r3)dr] =
A = Q/4pi*Integration [r2R-r3)dr]
sher mohammad
faculty askiitians
b.tech, iit delhi